RC Meeting, Friday, October 29 , 2021
RC Meeting Information
RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920
- Beam delivery to Halls B,C,D
Hall C Status
Continue production at 42 muA.
- day shift: productive, start BCM calibration ~14:10. D. Mack fixed the drifting ibcm1 by removing the 0.25V offset electronic module.
- swing shift: finish BCM calibration. resume production
- owl shift: very productive. dummy target run toward the morning. some pedestals need attention/updating (SHMS hodoscope 2Y, HMS calo D...).
Number of good pions is 6633 (not counting overnight harvest).
Current Run Plan
- Maintenance day shifted to 11/03/2021. Acceleration retuning firmly scheduled for 11/05/2021-11/07/2021. all downtime starts at or around 08:00 so people/groups that need access in the Hall might want to take advantage of these periods.
- Continue data taking until early Nov
- Given the uncertainty associated with recent Hall A events it is not clear that the maintenance and retuning will happen as scheduled (Tu and Th, respectively). Need to follow this up with the scheduling committee.
- Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
- HMS Q3 insulating vacuum spikes generating annoyance alarms that take minutes to recover. https://logbooks.jlab.org/entry/3930175 Noise levels on transducers look OK, but Junaid's plot indicates Q3 is burping: https://logbooks.jlab.org/entry/3932992 .
- Serious HB LHe low liquid level alarms. https://logbooks.jlab.org/entry/3933051 Always seems to start with a subtle, simultaneous increase in level in Q1/2/3. https://logbooks.jlab.org/entry/3929152 .
- Target ladder camera was identified to be bad: Joe Beaufait will order camera ( PENDING )
- Still running with no UPS on the SHMS cryo controls - Steve L. has ordered batteries and UPS. ( PENDING )
- HCCAM02 not functioning (lift required to replace).
- How the harp scans and carbon hole target should look (see links):
- carbon hole target beam centering procedure:
- Resetting the "ioc heartbeat epics communications" (start by clicking on the heart at upper left of the magnet controls screen
- Fall 2021 Shift Schedule
- Hall-C Live Status
- Hall-C Logbook
- MCC White Board